OpenStudy (grace111):

Please help will reward medal . An iced cake costs 10p more than a plain cake . Sarah bought two of each cake. They cost £1 altogether . What is the cost of an iced cake ? Show your working Thanks ^_^

OpenStudy (anonymous):

let x be iced cake and y be plain cake iced is 10p more than plain so x = y + 10 sum is 1 so x + y = 1 or y + 10 + y = 1 solve for y then you can solve for x

OpenStudy (grace111):

thanks :)